OpenStudy (anonymous):

Solve the equation. 0.25 r – 0.125 + 0.5 r = 0.5 + r

OpenStudy (anonymous):

0.25 r – 0.125 + 0.5 r = 0.5 + r First get all the r terms together like so 0.25r+0.5r-r=0.125-0.5 than solve for problem hopes this hels